OpenStudy (anonymous):

Hi :) Please help What is the suns position in the solar system? What is its size(KM), what is its mass (KG) and what is its temperature at core in degrees.

OpenStudy (anonymous):

At the center of the solar system.

OpenStudy (anonymous):

It's diameter is about \(1.4 \times 10 ^6 KM\). Mass is around \(1.99 \times 10^{30} Kg\) . Temperature at the core is \(1.57×10^7 K\) .

OpenStudy (anonymous):

Thanks, but i need the core temperature in °C

OpenStudy (anonymous):

DO U KNOW?? ILL BE VERY GRATEFUL :)

OpenStudy (anonymous):

You have to subtract 273 from the temperature in K. 273 is very small compared to 1.57×10^7 K. So, the temperature is \(1.57 \times 10^7 C\)

OpenStudy (anonymous):

IN DEGREES PLEASE °C :)

OpenStudy (anonymous):

DONT SUPPOSE YOULL KNOW THIS TOO: What percentage of earth is water? What percentage of earth is land? What elements are found in our air and atmosphere? THANKS :)

OpenStudy (anonymous):

THE SUN: Length of life to date: Estimated life remaining:

OpenStudy (anonymous):

What elements are found in our air and atmosphere?

OpenStudy (vincent-lyon.fr):

Just read: http://en.wikipedia.org/wiki/Sun